# On System Operators with Variation Bounding Properties

### v2

@@ 1 Introduction
$A\in\mathds{R}^{n\times n}$, $b,c^{\mathsf{T}}\in\mathds{R}^{n}$, that map nonnegative inputs $u$ to nonnegative outputs $y$ are characterized by nonnegative impulse responses $g(t):=cA^{t-1}b\geq 0$, $t\geq 1$ and referred to as externally positive [7]. In the particular case that $A$, $b$ and $c$ are element-wise nonnegative, the system is called internally positive as it is externally positive and preserves the positivity of the state for nonnegative inputs. Systems of this kind frequently occur through compartmental modelling [8, 2] and are known for several advantageous analytical properties such as scalable stability certificates [29], optimal controller design [33, 25] or the avoidance of over- and undershooting in closed-loop design [15, 4, 27, 32].
@@
In recent years, efforts have been made to study systems with extended positivity properties based on their influence on the variation, i.e., number of sign changes, of the input signal $u$ or the state $x(t)$ (see, e.g., [35, 21, 16, 14]. A mapping $u\mapsto Xu$ is called $k$-variation bounding ($\text{VB}_{k}$) if it maps an input $u$ with at most $k$ sign changes to an output $Xu$ of at most $k$ sign changes. If the mapping is $\text{VB}_{j}$ for all $j\in\{0,\dots,k\}$, it is called $k$-variation diminishing ($\text{VD}_{k}$) and if, additionally, the order of sign changes is preserved whenever $u$ and $Xu$ have equal variation, the $\text{VD}_{k}$ property is called order-preserving ($\text{OVD}_{k}$).
@@
Systems with input-output $\text{OVD}_{k}$ properties have been studied in terms of Hankel and Toeplitz operator [16], which resulted in intermediates between external positivity ($\text{OVD}_{0}$ operators) and the well-known class of relaxation systems [37] ($\text{OVD}_{\infty}$ Hankel operators). These intermediates have been exploited in [12] to provide guarantees for internal positivity preserving model order reduction – a property well-known for $\text{OVD}_{\infty}$ Hankel operators. Studies of autonomous (unforced) systems with $\text{VD}_{k}$ and $\text{VB}_{k}$ state-transition maps [21, 36, 23] have provided an extended, unified approach to several important non-linear system analysis tools such as monotonicity/cooperativity, contraction analysis or Poincare-Bendixson results. A system class that was proposed to merge the input-output and autonomous case has been suggested in [14]. This resulted in a first certificate for $\text{OVD}_{k}$ observability/controllability operators and its application to bounding the variation of an impulse response beyond external positivity. While the latter problem is of interest, e.g., when bounding the number of over- and undershoots in a step response or the number of bangs in bounded optimal control problems [20, 10], it is still an open problem, where several lower bounds [3, 31, 5] have been derived, but only few upper bounds [5].
@@
The goal of this work is to derive a tractable characterization of systems with $\text{VB}_{k}$ observability/controllability operators. This complements the literature in several aspects. Firstly, we close the gap of algebrapenic characterization of $\text{VB}_{k}$ and $\text{VD}_{k}$ matrices beyond limitations to rank and dimensions (cf. [35, 19]) or $\text{OVD}_{k}$ matrices (cf. [19, 16]). Based on that, computationally tractable certificates are achieved by extending related results in [26]. Secondly, it is shown that an application of these results to the infinite-dimensional observability/controllability operator leads to the external positivity of a related family of linear time-invariant systems. This is similar to the approach in [16], but goes beyond $\text{OVD}_{k}$ and does not requiring $A$ to be $\text{OVD}_{k}$ (cf. [14]). To underline this benefit, we present a system example for which there exists a realization with $\text{VD}_{2}$ observability operator, but no realization with an $\text{VD}_{2}$ AA. This mimics the fact that not all externally positive systems have an internally positive realization [1, 7, 15]. Further, as any Hankel operator factors into observability and controllability operator, our results also provide a first sufficient certificate of $\text{VB}_{k}$ and $\text{VD}_{k}$ Hankel operators. Our results are illustrated by bounding the variation of the impulse response aka. the number of over- and undershoots in an step response. In future work, we hope to use these insights to also derive characteristics in terms of zeros and poles. Lastly, the presented results have been instrumental to subsequent works dealing with sparse optimization problems [22] and self-oscillations in relay feedback systems [13, 34].
@@
Finally note that this paper is an extension of the authors’ earlier work [30]. The present work differs from [30] in several major aspects: (i) previously only the special case of strict variation bounding was considered, which could be built upon existing matrix theory results [26, 19]. As mentioned earlier, the current manuscript extend the system operator and matrix theory results substantially to the non-strict case. In particular, much weaker assumptions on rank and dimensions are made compared to [35, 19]. These novel results are valuable independent of the considered application and also directly include the case of (non-order-persevering) $k$-variation diminishment; (ii) the current paper includes an additional example to illustrate further benefits and insights compared to existing work [14]; (iii) our earlier results on strict variation bounding have been published without proofs. These proofs can only be found in the present manuscript.
@@ 2.1 Notations
We write $\mathds{Z}$ for the set of integers and $\mathds{R}$ for the set of reals, with $\mathds{Z}_{\geq 0}$ and $\mathds{R}_{\geq 0}$ standing for the respective subsets of nonnegative elements – the corresponding notations are also used for subsets starting from non-zero values, strict inequality as well as reversed inequality signs. The set of real sequences with indices in $\mathds{Z}$ is denoted by $\mathds{R}^{\mathds{Z}}$. For matrices $X=(x_{ij})\in\mathds{R}^{n\times m}$, we say that $X$ is nonnegative, $X\geq 0$ or $X\in\mathds{R}^{n\times m}_{\geq 0}$ if all elements $x_{ij}\in\mathds{R}_{\geq 0}$ – corresponding notations are used for matrices with strictly positive entries and reversed inequality signs. These notations are also used for sequences $x=(x_{i})\in\mathds{R}^{\mathds{Z}}$. For $k,l\in\mathds{Z}$, we write $(k:l):=\{k,k+1,\dots,l\}$, $k\leq l$. In the case $k>l$, the notation represents the empty set. The complement of a $k$-tuple, II, is ${I}^{\text{c}}$ such that $I\cup{I}^{\text{c}}=(1:n)$. If $X\in\mathds{R}^{n\times n}$, then $\sigma(X)=\{\lambda_{1}(X),\dots,\lambda_{n}(X)\}$ denotes its spectrum, where the eigenvalues are ordered by descending absolute value, i.e., $\lambda_{1}(X)$ is the eigenvalue with the largest magnitude, counting multiplicity. If the magnitude of two eigenvalues coincides, we sub-sort them by decreasing real part. The identity matrix in $\mathds{R}^{n\times n}$ is denoted by $I_{n}$ or if the dimensions are obvious, simply by $I$. A (consecutive) $j$-minor of $X\in\mathds{R}^{n\times m}$ is a minor which is constructed of $j$ columns and $j$ rows of $X$ (with consecutive indices). The set of $j$-consecutive minors with column (row) indices $(1:j)$ are said to be the $j$-column(row) initial minors. The submatrix with rows $I\subset(1:n)$ and columns $J\subset(1:m)$ is written as ${X}_{I,J}$. In case of subvectors, we simply write $x_{I}$. With slight abuse of notation, we also use this to denote subsets of ordered index sets $x\in\mathcal{I}_{n,r}$, where
$\mathcal{I}_{n,r}:=\{v=\{v_{1},\dots,v_{r}\}\subset\mathds{N}:1\leq v_{1}<\dots<v_{r}\leq n\}.$
Finally, if $X\in\mathds{R}^{n\times n}$ is such that $\det{X}_{(2:n-1),(2:n-1)}\neq 0$, then the following identity exists [18, Equation 0.8.11]:
$\det{X}_{(1:n),(1:n)}\det{X}_{(2:n-1),(2:n-1)}\\ =\det{X}_{(1:n-1),(1:n-1)}\det{X}_{(2:n),(2:n)}-\det{X}_{(1:n-1),(2:n)}\det{X}_{(2:n),(1:n-1)}$ (2)
@@ 2.2 Variation diminishing maps
where $\bar{u}$ is the vector resulting from replacing zeros by elements that maximize the resulting sum. Obviously, $\text{S}^{-}(u)\leq S^{+}(u)$, but equality does not necessarily need to hold, e.g., $\text{S}^{-}(\begin{pmatrix}1&0&2\end{pmatrix})=\text{S}^{-}(\begin{pmatrix}1&2\end{pmatrix})=0$, but $S^{+}(\begin{pmatrix}1&0&2\end{pmatrix})=\text{S}^{-}(\begin{pmatrix}1&-1&2\end{pmatrix})=2$.
@@ 2.3 Total Positivity Theory
Total positivity theory [19] is known to provide algebraic conditions for the $\text{OVD}_{k-1}$ and $\text{SVB}_{k-1}$ property by means of compound matrices. For $X\in\mathds{R}^{n\times m}$, the $(i,j)$-th entry of the so-called r-th multiplicative compound matrix $X_{[r]}\in\mathds{R}^{\binom{n}{r}\times\binom{m}{r}}$ is defined via $\det({X}_{I,J})$, where $I$ and $J$ are the $i$-th and $j$-th element of the $r$-tuples in $\mathcal{I}_{n,r}$ and $\mathcal{I}_{m,r}$, respectively, assuming lexicographical ordering. For example, if $X\in\mathds{R}^{3\times 3}$, then
@@
equals $X_{[2]}$. In our derivations, the following properties of the multiplicative compound matrix will be elementary (see, e.g., [9, Section 6] and [17, Subsection 0.8.1]).
@@
It is readily seen that $X$ is $\text{OVD}_{0}$ if only if $X=X_{[1]}\geq 0$. This equivalence can be generalized to higher orders by the following definitions and characterizations (see [16, Prop. 7] and [19, Theorem 5.1.1]).
@@
Note that one is only interested in cases with $n>m$, because any $X\in\mathds{R}^{n\times n}$ is $\text{SVB}_{n-1}$. In order to extend strict to the non-strict cases, the following proposition (see, e.g., [19, Propositions 5.1.1 & 5.1.2], [14, Propositions 2.7]) and lemma [19, Lemma 5.1.1] will be used.
@@
For example, Proposition 3 in conjunction with Lemma 2 characterizes $\text{VB}_{m-1}$ by ${SC}_{m}$ as in [19, Theorems 5.1.3 & 5.1.3’].
@@
Next, we will review how checking the sign of the elements in $X_{[k]}$ can be simplified compared to computing all of its entries. This does will not only make $\text{(S)}\text{VB}_{k-1}$ more tractable, but is particularly important for our applications to infinite-dimensional operators. We begin with a sufficient condition for $m$-sign consistency [19, Theorem 2.3.1], which only requires to consider consecutive row and column initial minors.
@@
An application of this result yields the following (sufficient) certificate (see [16, Proposition 8]) for (non-)strict $k$-positivity.
@@
A related results is the following sufficient and necessary condition for ${SSC}_{m}$ [26, Theorem 2.2] and its application to ${SSC}_{k}$ case.
@@
Proposition 7 utilizes a bijection between the $m$-order minors of a matrix $X\in\mathds{R}^{n\times m}$ and the minors of a related transformed matrix [26], which we will also use for a new non-strict counter-parts.
@@ 2.4 Variation diminishing observability operators
As discussed in [14, Lemma 3.4], the $\text{OVD}_{k}$ property of ${\mathcal{O}}(A,c)$ is equivalent to
@@
being $\text{OVD}_{k}$ for all $t\geq k$. For brevity, we will mostly drop $(A,c)$ and write $\mathcal{O}$} and $\mathcal{O}^{t}$}, respectively. A sufficient condition based on the assumption of a $k$-positive AA was shown in [14, Theorem 3.5].
@@ 3.1 The $\text{SVB}_{k-1}$ case
Analogously to the $\text{OVD}_{k-1}$ case, one can show that $\mathcal{O}$} is $\text{SVB}_{k-1}$ if and only if $\mathcal{O}^{t}$} is $\text{SVB}_{k-1}$ for all $t\geq k-1$. Thus, by Proposition 2, it suffices to check ${SSC}_{n-1}$ of $\mathcal{O}^{t}$} for all $t\geq n-1$. Using Proposition 7, this is equivalent to the strict positivity/negativity of certain sequences of n1n-1-minors. Next, we will show that these sequences of n1n-1- minors correspond to impulse responses of related LTI system. Checking $\text{SVB}_{k-1}$ then becomes equivalent to the verification of strict external positivity of $n$ LTI systems for which numerically efficient certificates can be found, e.g., in [15, 32].
@@
In order to proceed with the general $\text{SVB}_{k-1}$ case, we need the following generalization of Proposition 2, which has appeared earlier in several related forms (see, e.g., [19, 35]), but to the best of our knowledge not been stated and derived in given full generality.
@@
Since strict externally positivity requires a dominant positive pole (see, e.g., [24, 8, 1]), the following eigenvalue constraint can be shown as a consequence of Item ii in Lemma 1.
@@
Let observable $(A,c)$ be such that $\mathcal{O}$ is ${SSC}_{k}$. Then, $\lambda_{1}(A),\dots,\lambda_{k}(A)$ are simple and strictly positive.
Interestingly, this necessary condition is also implied by $k$-positivity of $A$ as in Proposition 8 as well as by the $k$-positivity of the Hankel and Toeplitz operators [16]. In particular, if $k=n$, then there exists a $T\in\mathds{R}^{n\times n}$ such that $T^{-1}AT=\textnormal{diag}(\lambda_{1}(A),\dots,\lambda_{n}(A))\geq 0$ and $cT=\begin{pmatrix}1&\cdots&1\end{pmatrix}$, which as shown in [14] fulfills the requirements of Proposition 8.

### v3

@@ 1 Introduction
$A\in\mathds{R}^{n\times n}$, $b,c^{\mathsf{T}}\in\mathds{R}^{n}$, that map nonnegative inputs $u$ to nonnegative outputs $y$ are characterized by nonnegative impulse responses $g(t):=cA^{t-1}b\geq 0$, $t\geq 1$ and referred to as externally positive [farina2011positive]. In the particular case that $A$, $b$ and $c$ are element-wise nonnegative, the system is called internally positive as it is externally positive and preserves the positivity of the state for nonnegative inputs. Systems of this kind frequently occur through compartmental modelling [brown1980compartmental, farina2011positive] and are known for several advantageous analytical properties such as scalable stability certificates [rantzer2015scalable], optimal controller design [tanaka2011bounded, pates2019optimal] or the avoidance of over- and undershooting in closed-loop design [grussler2019tractable, darbha2003synthesis, phillips1988conditions, taghavian2023external].
@@
In recent years, efforts have been made to study systems with extended positivity properties based on their influence on the variation, i.e., number of sign changes, of the input signal $u$ or the state $x(t)$ (see, e.g., [weiss2019generalization, margaliot2018revisiting, grussler2020variation, grussler2021internally]. A mapping $u\mapsto Xu$ is called $k$-variation bounding ($\text{VB}_{k}$) if it maps an input $u$ with at most $k$ sign changes to an output $Xu$ of at most $k$ sign changes. If the mapping is $\text{VB}_{j}$ for all $j\in\{0,\dots,k\}$, it is called $k$-variation diminishing ($\text{VD}_{k}$) and if, additionally, the order of sign changes is preserved whenever $u$ and $Xu$ have equal variation, the $\text{VD}_{k}$ property is called order-preserving ($\text{OVD}_{k}$).
@@
Systems with input-output $\text{OVD}_{k}$ properties have been studied in terms of Hankel and Toeplitz operator [grussler2020variation], which resulted in intermediates between external positivity ($\text{OVD}_{0}$ operators) and the well-known class of relaxation systems [willems1976realization] ($\text{OVD}_{\infty}$ Hankel operators). These intermediates have been exploited in [grussler2020balanced] to provide guarantees for internal positivity preserving model order reduction – a property well-known for $\text{OVD}_{\infty}$ Hankel operators. Studies of autonomous (unforced) systems with $\text{VD}_{k}$ and $\text{VB}_{k}$ state-transition maps [margaliot2018revisiting, weiss2021generalization, ofir2024contraction] have provided an extended, unified approach to several important non-linear system analysis tools such as monotonicity/cooperativity, contraction analysis or Poincaré-Bendixson results. A system class that was proposed to merge the input-output and autonomous case has been suggested in [grussler2021internally]. This resulted in a first certificate for $\text{OVD}_{k}$ observability/controllability operators and its application to bounding the variation of an impulse response beyond external positivity. While the latter problem is of interest, e.g., when bounding the number of over- and undershoots in a step response or the number of bangs in bounded optimal control problems [luenberger1968optimization, friesz2010dynamic], it is still an open problem, where several lower bounds [damm2014zero, swaroop1996some, elkhoury1993discrete] have been derived, but only few upper bounds [elkhoury1993discrete].
@@
The goal of this work is to derive a tractable characterization of systems with $\text{VB}_{k}$ observability/controllability operators, which complements the literature in several aspects: firstly, we close the gap of algebraic characterization of $\text{VB}_{k}$ and $\text{VD}_{k}$ matrices beyond limitations to rank and dimensions (cf. [weiss2019generalization, karlin1968total]) or $\text{OVD}_{k}$ matrices (cf. [karlin1968total, grussler2020variation]). Based on that, computationally tractable certificates are achieved by extending related results in [pena_matrices_1995]. Secondly, it is shown that an application of these results to the infinite-dimensional observability/controllability operator leads to the external positivity of a related family of linear time-invariant systems. This is similar to the approach in [grussler2020variation], but goes beyond $\text{OVD}_{k}$ and does not require $A$ to be $\text{OVD}_{k}$ (cf. [grussler2021internally]). To underline this benefit, we present a system example for which there exists a realization with $\text{VD}_{2}$ observability operator, but no realization with an $\text{VD}_{2}$ AA. This parallels the fact that not all externally positive systems have an internally positive realization [benvenuti2004tutorial, farina2011positive, grussler2019tractable]. Further, as any Hankel operator factors into observability and controllability operator, our results also provide a first sufficient certificate of $\text{VB}_{k}$ and $\text{VD}_{k}$ Hankel operators. Our results are illustrated by bounding the variation of the impulse response aka. the number of over- and undershoots in a step response. In future work, we hope to use these insights to also derive characteristics in terms of zeros and poles. Lastly, the presented results have been instrumental to subsequent works dealing with sparse optimization problems [marmary2025tractable] and self-oscillations in relay feedback systems [grussler2025discrete, tong2025selfsustained].
@@
Finally note that this paper is an extension of the authors’ earlier work [roth2024variation]. The present work differs from [roth2024variation] in several major aspects: (i) previously only the special case of strict variation bounding was considered, which could be built upon existing matrix theory results [pena_matrices_1995, karlin1968total]. As mentioned earlier, the current manuscript extends the system operator and matrix theory results substantially to the non-strict case. In particular, much weaker assumptions on rank and dimensions are made compared to [weiss2019generalization, karlin1968total]. These novel results are valuable independent of the considered application and also directly include the case of (non-order-preserving) $k$-variation diminishment; (ii) the current paper includes an additional example to illustrate further benefits and insights compared to existing work [grussler2021internally]; (iii) our earlier results on strict variation bounding have been published without proofs. These proofs can only be found in the present manuscript.
@@ 2.1 Notations
We write $\mathds{Z}$ for the set of integers and $\mathds{R}$ for the set of reals, with $\mathds{Z}_{\geq 0}$ and $\mathds{R}_{\geq 0}$ standing for the respective subsets of nonnegative elements – the corresponding notations are also used for subsets starting from non-zero values, strict inequality as well as reversed inequality signs. The set of real sequences with indices in $\mathds{Z}$ is denoted by $\mathds{R}^{\mathds{Z}}$. For matrices $X=(x_{ij})\in\mathds{R}^{n\times m}$, we say that $X$ is nonnegative, $X\geq 0$ or $X\in\mathds{R}^{n\times m}_{\geq 0}$ if all elements $x_{ij}\in\mathds{R}_{\geq 0}$ – corresponding notations are used for matrices with strictly positive entries and reversed inequality signs. These notations are also used for sequences $x=(x_{i})\in\mathds{R}^{\mathds{Z}}$. For $k,l\in\mathds{Z}$, we write $(k:l):=\{k,k+1,\dots,l\}$, $k\leq l$. In the case $k>l$, the notation represents the empty set. If $X\in\mathds{R}^{n\times n}$, then $\sigma(X)=\{\lambda_{1}(X),\dots,\lambda_{n}(X)\}$ denotes its spectrum, where the eigenvalues are ordered by descending absolute value, i.e., $\lambda_{1}(X)$ is the eigenvalue with the largest magnitude, counting multiplicity. If the magnitude of two eigenvalues coincides, we sub-sort them by decreasing real part. The identity matrix in $\mathds{R}^{n\times n}$ is denoted by $I_{n}$ or if the dimensions are obvious, simply by $I$. A (consecutive) $j$-minor of $X\in\mathds{R}^{n\times m}$ is a minor which is constructed of $j$ columns and $j$ rows of $X$ (with consecutive indices). The set of $j$-consecutive minors with column (row) indices $(1:j)$ are said to be the $j$-column(row) initial minors. The submatrix with rows $I\subset(1:n)$ and columns $J\subset(1:m)$ is written as ${X}_{I,J}$. In case of subvectors, we simply write $x_{I}$. With slight abuse of notation, we also use this to denote subsets of ordered index sets $x\in\mathcal{I}_{n,r}$, where
$\mathcal{I}_{n,r}:=\{v=\{v_{1},\dots,v_{r}\}\subset\mathds{N}:1\leq v_{1}<\dots<v_{r}\leq n\},$
and the complement of $x$ is defined as the unique ${x}^{\text{c}}\in\mathcal{I}_{n,n-r}$ such that ${x}^{\text{c}}\cup x=(1:n)$. Finally, if $X\in\mathds{R}^{n\times n}$ is such that $\det{X}_{(2:n-1),(2:n-1)}\neq 0$, then the following identity exists [horn2012matrix, Equation 0.8.11]:
$\det{X}_{(1:n),(1:n)}\det{X}_{(2:n-1),(2:n-1)}\\ =\det{X}_{(1:n-1),(1:n-1)}\det{X}_{(2:n),(2:n)}\\ -\det{X}_{(1:n-1),(2:n)}\det{X}_{(2:n),(1:n-1)}$ (2)
@@ 2.2 Variation diminishing maps
[…]
@@ 2.3 Total Positivity Theory
Total positivity theory [karlin1968total] is known to provide algebraic conditions for the $\text{OVD}_{k-1}$ and $\text{SVB}_{k-1}$ property by means of compound matrices. For $X\in\mathds{R}^{n\times m}$, the $(i,j)$-th entry of the so-called r-th multiplicative compound matrix $X_{[r]}\in\mathds{R}^{\binom{n}{r}\times\binom{m}{r}}$ is defined via $\det({X}_{I,J})$, where $I$ and $J$ are the $i$-th and $j$-th element of the $r$-tuples in $\mathcal{I}_{n,r}$ and $\mathcal{I}_{m,r}$, respectively, assuming lexicographical ordering. For example, if $X\in\mathds{R}^{3\times 3}$, then
@@
equals $X_{[2]}$. In our derivations, the following properties of the multiplicative compound matrix will be elementary (see, e.g., [fiedler2008special, Section 6] and [horn2012matrix, Subsection 0.8.1]).
@@
It is readily seen that $X$ is $\text{OVD}_{0}$ if only if $X=X_{[1]}\geq 0$. This equivalence can be generalized to higher orders by the following definitions and characterizations (see [grussler2020variation, Prop. 7] and [karlin1968total, Theorem 5.1.1]).
@@
Note that one is only interested in cases with $n>m$, because any $X\in\mathds{R}^{n\times n}$ is $\text{SVB}_{n-1}$. In order to extend strict to the non-strict cases, the following proposition (see, e.g., [karlin1968total, Propositions 5.1.1 & 5.1.2], [grussler2021internally, Propositions 2.7]) and lemma [karlin1968total, Lemma 5.1.1] will be used.
@@
For example, Proposition 3 in conjunction with Lemma 2 characterizes $\text{VB}_{m-1}$ by ${SC}_{m}$ as in [karlin1968total, Theorems 5.1.3 & 5.1.3’].
@@
Next, we will review how checking the sign of the elements in $X_{[k]}$ can be simplified compared to computing all of its entries. This does will not only make $\text{(S)}\text{VB}_{k-1}$ more tractable, but is particularly important for our applications to infinite-dimensional operators. We begin with a sufficient condition for $m$-sign consistency [karlin1968total, Theorem 2.3.1], which only requires to consider consecutive row and column initial minors.
@@
An application of this result yields the following (sufficient) certificate (see [grussler2020variation, Proposition 8]) for (non-)strict $k$-positivity.
@@
A related result is the following sufficient and necessary condition for ${SSC}_{m}$ [pena_matrices_1995, Theorem 2.2] and its application to ${SSC}_{k}$ case.
@@
Proposition 7 utilizes a bijection between the $m$-order minors of a matrix $X\in\mathds{R}^{n\times m}$ and the minors of a related transformed matrix [pena_matrices_1995], which we will also use for a new non-strict counter-parts.
@@ 2.4 Variation diminishing observability operators
As discussed in [grussler2021internally, Lemma 3.4], the $\text{OVD}_{k}$ property of ${\mathcal{O}}(A,c)$ is equivalent to
@@
being $\text{OVD}_{k}$ for all $t\geq k$. For brevity, we will mostly drop $(A,c)$ and write $\mathcal{O}$} and $\mathcal{O}^{t}$}, respectively. A sufficient condition based on the assumption of a $k$-positive AA was shown in [grussler2021internally, Theorem 3.5].
@@ 3.1 The $\text{SVB}_{k-1}$ case
Analogously to the $\text{OVD}_{k-1}$ case, one can show that $\mathcal{O}$} is $\text{SVB}_{k-1}$ if and only if $\mathcal{O}^{t}$} is $\text{SVB}_{k-1}$ for all $t\geq k-1$. Thus, by Proposition 2, it suffices to check ${SSC}_{n-1}$ of $\mathcal{O}^{t}$} for all $t\geq n-1$. Using Proposition 7, this is equivalent to the strict positivity/negativity of certain sequences of n1n-1-minors. Next, we will show that these sequences of n1n-1- minors correspond to impulse responses of related LTI system. Checking $\text{SVB}_{k-1}$ then becomes equivalent to the verification of strict external positivity of $n$ LTI systems for which numerically efficient certificates can be found, e.g., in [grussler2019tractable, taghavian2023external].
@@
In order to proceed with the general $\text{SVB}_{k-1}$ case, we need the following generalization of Proposition 2, which has appeared earlier in several related forms (see, e.g., [karlin1968total, weiss2019generalization]), but to the best of our knowledge not been stated and derived in given full generality.
@@
Since strict externally positivity requires a dominant positive pole (see, e.g., [ohta1984reachability, farina2011positive, benvenuti2004tutorial]), the following eigenvalue constraint can be shown as a consequence of Item ii in Lemma 1.
@@
Let $(A,c)$ be observable such that $\mathcal{O}$ is ${SSC}_{k}$. Then, $\lambda_{1}(A),\dots,\lambda_{k}(A)$ are simple and strictly positive.
Interestingly, this necessary condition is also implied by $k$-positivity of $A$ as in Proposition 8 as well as by the $k$-positivity of the Hankel and Toeplitz operators [grussler2020variation]. In particular, if $k=n$, then there exists a $T\in\mathds{R}^{n\times n}$ such that $T^{-1}AT=\textnormal{diag}(\lambda_{1}(A),\dots,\lambda_{n}(A))\geq 0$ and $cT=\begin{pmatrix}1&\cdots&1\end{pmatrix}$, which as shown in [grussler2021internally] fulfills the requirements of Proposition 8.